40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5244     Entry time: Tue Aug 16 04:25:34 2011
Author: Suresh, Kiwamu 
Type: Update 
Category: SUS 
Subject: alignment of MC output to Y-arm using PZTs 

We did several things today+night.  The final goal was to lock the PRC so that we could obtain the POX, POY and POP beams.  However there were large number of steps to get there.

1) We moved the ITMY into its place and balanced the table

2) We then aligned the Y-arm cavity to the green beam which was set up as a reference before we moved the ETMY and ITMY to adjust the OSEMS.  We had the green flashing in Y-arm

3) We checked the beam position on PR2. It was okay. This confirmed that we were ready to send the beam onto the Y arm.

4) We then roughly aligned the IR beam on ETMY where Jamie had placed an Al foil with a hole.  We got the arm flashing in both IR and green. 

5) We used the PZTs to make the green and IR beams co-incident and flashing in the Y arm.  This completed the alignment of the IR beam into the Y-arm.

6) The IPPO (pick-off) window had to be repositioned to avoid clipping.  The IPANG beam was aligned such that it exits the ETMY chamber onto the ETMY table.  It can now be easily sent to the IPANG QPD.

7) Then BS was aligned to direct the IR beam into the X-arm and had the X-arm flashing.  It had already been aligned to its green.

8) It was now the turn of the SRC.  The beam spots on all the SRC related optics were off centered.  We aligned all the optics in the AS path to get the AS beam on to the AP table.

9) The AS beam was very faint so we repositioned the AS camera to the place intended for AS11 PD, since there was a brighter beam available there. 

10) We could then obtain reflections from ITMY, ITMX and PRM at the AS camera. 

11) Problems:

      a) ITMY osems need to be readjusted to make sure that they are in mid-range.  Several are out of range and so the damping is not effective.

      b) When we tried to align SRC the yaw OSEM had to be pushed to its full range.  We therefore have to turn the SRM tower to get it back into range.

 12)  We stopped here since moving the SRM is not something to be attempted at the end of a rather long day. Kiwamu is posting a plan for the rest of the day.

ELOG V3.1.3-